Q2: Marres limited has the following demand and cost functions



Demand function: P=80-3Q, where P is the unit selling price and Q is quantity in thousands



Cost function: TC=Q2+20Q+100, where TC is total cost in Ksh 000000



Required:



Optimal price to maximize profit (3mks)



Maximum profit (2mks)

1
Expert's answer
2022-03-02T17:04:04-0500

Given that:

"TC=Q^2+20Q+100" and

"P=80-3Q"

"TR=P\u00d7Q"

"TR=(80-3Q)Q=80Q-3Q^2"

"TR'=MR=80-6Q"

"TC'=MC=2Q+20"


"MR=MC" at equilibrium

Then:

"80-6Q=2Q+20"

"8Q=60"

"Q^*=7.5" in thousand


a) Optimal price to maximize profit:

"P^*=80-3(7.5)=Ksh57.5"


b) Maximum Profit, "\pi" ="TR-TC"

At "Q^*=7.5"

"\pi=80(7.5)-3(7.5)^2-((7.5)^2+20(7.5)+100)"

"\pi=125(Ksh '000000)"